OpenStudy (anonymous):

I have the problem 27^4z=9^(z+1) I don't want the answer, just help figuring out the proper way to simplify both sides of the equation so I can add the exponents. I know the bases will be the same, 3^(whatever power)

OpenStudy (anonymous):

I have gotten as far as 3^3(4z)=3^2(z+1)

OpenStudy (amistre64):

log both sides

OpenStudy (anonymous):

I don't know what to do with the exponents, whether I should add them, multiply them, divide them or whatever else. I am not familiar with the log function yet. The chapter the homework is in is on exponential functions, so trying to do it that way first.

OpenStudy (amistre64):

27^4z=9^(z+1) log(27^4z) = log(9^(z+1)) 4z (log(27)) = (z+1) (log(9)) get like terms together now

OpenStudy (anonymous):

I already have the bases, I just need to know what to do with the exponents. Again, trying to do this without a calculator. I want to understand the process before I rely on a calculator for the answers

OpenStudy (amistre64):

exponents an logs are hand in hand. they are inverses of each other and you need logs to work exponents with a variable in them

OpenStudy (amistre64):

you can "remember" exponents if you want and try to recall what is what... but it is harder

OpenStudy (anonymous):

4z (log(27)) = (z+1) (log(9)) 4z.3log (3)= (z+1)2log (3) 12z=2z+2 z=1/5

OpenStudy (amistre64):

the relation is this: B^x = y x = logB(y)

OpenStudy (amistre64):

3^2 = 9 2 = log3(9)

OpenStudy (amistre64):

As for exponent rules: 3^(2^x) = 3^(2x). when exponents have an exponent; the exponents multiply together

OpenStudy (anonymous):

THAT is what I needed :D The exponent rule.

OpenStudy (anonymous):

I got it right easily that time. Just didn't know what to do with it, thank you very much d:D

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

{I have gotten as far as 3^3(4z)=3^2(z+1)}. when you got this , why are you mess with log.. just compare exponents of 3 n get the result

OpenStudy (amistre64):

easy when you can "know" the base..